OpenStudy (anonymous):

Flint is purchasing a home for $115,750 and is financing 85% of it. The documentary stamp tax on the deed in his state is $0.70 per $100 portion thereof. How much will he pay for the documentary stamp tax? a.$1405.54 b.$1653.57 c.688.71 d.810.25

OpenStudy (anonymous):

Please help! I don't understand any of this at all and I am a little behind and have to have ALL of my school stuff finished no later that may 12th or I don't graduate.

OpenStudy (anonymous):

@undeadknight26 @maryfitzgerald14 @swaggd2 @jeffyblood @JennyK1 @Chloeluvsyah

OpenStudy (anonymous):

@satellite73 @sammixboo

OpenStudy (anonymous):

@Hero @mathstudent55 @mathmate

OpenStudy (anonymous):

@ganeshie8 @Kainui

OpenStudy (anonymous):

I don't understand any of this at all :(

OpenStudy (anonymous):

First, lets find how much flint is financing. Flint is financing 85% of 115, 750 dollars, which is: 0.85 * 115750 = 98387.5 dollars. Now the rate of this 'tax stamp', do you know what that is?

OpenStudy (anonymous):

no I don't.

OpenStudy (anonymous):

It's $0.70 per $100, which can be described as 0.7/100 which is just 0.007. Does that make sense?

OpenStudy (anonymous):

Yes it does.

OpenStudy (anonymous):

So what a tax is, is just a percentage of a certain total. This is how a tax is calculated. So if we know the tax 'rate' and some amount (i.e how much flint is financing), we know that the total tax flint has to pay is: Tax = How much flint is financing * tax rate. Can you figure out the answer based on that?

OpenStudy (anonymous):

So would it be .085(115,750)*.007?

OpenStudy (anonymous):

Yes. or more simply put: 98387.5 * 0.007

OpenStudy (anonymous):

So the answer is 688.71? C

OpenStudy (anonymous):

Would you be able to help me on another?

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

And I can try, please send me the link to the question.

OpenStudy (anonymous):

Jakob is purchasing a condominium for $347,500 and has financed 90% of it. He has purchased 2 discount points, has to pay a 1.75% origination fee of $100 plus .25% of the loan. What are his total fees that he must pay?

OpenStudy (anonymous):

Jakob has financed 347500 * 0.9 of a condo, which means Jakob has financed $312,750. This is our starting point.

OpenStudy (anonymous):

Okay, so from what I think I understand, he has to pay a fee of 100 dollars PLUS 0.25% of the loan right?

OpenStudy (anonymous):

ummm I think so

OpenStudy (anonymous):

Okay. So we need to figure out how much the loan was Jakob took out, which is simply 347500 - 312750 = $34,750 From here, we know he has to pay 0.25% of this loan back. So what is 0.25% of the $34,750?

OpenStudy (anonymous):

Hint, 0.25% is 0.0025 in decimal value.

OpenStudy (anonymous):

868.75

OpenStudy (anonymous):

wait 86.87

OpenStudy (anonymous):

Correct. So isn't the total loan amount he has to pay = 100 + 86.87? This question sure is worded strangely.

OpenStudy (anonymous):

hmmm? my answer choices are a.$12,610.01 B$13,031.25 c.$13,218.13 d.$13,913.13

OpenStudy (anonymous):

It's something to do with the discount points, I misinterpreted the question. Sorry but I can't see based on the explanation as to what a discount point even is, try asking the question in a seperate thread and maybe someone can help you. :)

OpenStudy (anonymous):

I know that 2 discount points equal .02 and that is supposed to be multiplied by the principle or something.

OpenStudy (anonymous):

okay, so I think it's B. Here's my calculation for B. 0.02 (Discount points) * 347500 = 6950 0.0175 (organization fee) * 347500 = 6081.25 6950 + 6081.25 = 13031.25

OpenStudy (anonymous):

ok thank you very much!!! your awesome Fzxt!